# Dahyeon Kye, Changhyun Roh, Sukhun Ko, Chanho Eom, and Jihyong Oh https://github.com/CMLab-Korea/Awesome-Video-Frame-Interpolation

### v1

@@ 2.2.3 Kernel- and Flow-based Combined
Kernel- and flow-based approaches each offer distinct strengths in VFI. Flow-based methods [16, 30, 7, 33, 74, 35, 37, 38, 97] estimate dense motion fields to align frames in a temporally coherent manner, but their performance degrades due to inaccurate flow estimation or the presence of occlusions. In contrast, kernel-based methods [28, 29, 16, 7, 31, 32, 33, 74, 34, 35, 36, 37, 8, 38, 39, 40, 41, 42, 97, 43, 44, 46, 56, 47] directly synthesize pixels using learned, spatially adaptive convolutional kernels, offering greater robustness in regions with complex motion. However, they are limited by their local receptive field and thus struggle with large displacements.
@@
Hybrid methods combine these complementary approaches by using flow estimation to guide the placement and orientation of learned kernels, yielding both global motion alignment and localized refinement. This combined approach [16, 30, 7, 33, 74, 35, 36, 37, 38, 97] typically begins by estimating optical flows using dedicated flow networks. Some methods adopt off-the-shelf optical flow networks [126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136] to guide the sampling location or trajectory of adaptive kernels. The kernels are then applied along flow-aligned paths to aggregate motion-aware pixel neighborhoods. MEMC-Net [33] exemplifies this design by integrating PWC-Net for flow estimation and deformable convolution [122] for localized refinement. In this setup, flow fields define the sampling offsets, while the kernel weights are learned to capture residual motion and restore high-frequency content. The predicted flows determine the sampling offsets for each pixel, while the learnable kernels capture residual motion and texture details. More recently, LADDER [47] introduces a lightweight encoder-decoder architecture that jointly estimates motion-aware features and spatially adaptive kernels, reducing complexity while maintaining hybrid modeling capacity.
@@ 2.2.4 Phase-based
An alternative direction in VFI explores the use of phase information to implicitly capture motion cues. In the frequency domain, pixel-wise representations can be decomposed into amplitude and phase components, where temporal phase shifts across frames encode the apparent motion of underlying structures. Phase-based methods [48, 49] exploit this property by estimating motion through local phase variations, rather than relying on explicit correspondence or pixel displacement. To extract and manipulate phase information, most methods adopt multi-scale frequency representations such as complex steerable pyramids [142, 143, 144]. Within this framework, motion is modeled by interpolating both phase and amplitude at each pyramid level. Meyer et al[48] solves this optimization problem explicitly, while later method PhaseNet [49] adopts end-to-end learning strategies.
@@
However, the effectiveness of these methods is fundamentally constrained by the assumption that motion can be approximated as local phase shifts. While this holds for small or moderate motion magnitudes, the assumption breaks down in the presence of large motion, leading to phase ambiguity and aliasing artifacts [145, 146]. As a result, phase-based methods often struggle to preserve fine spatial details or sharp boundaries in high-speed motion scenarios, limiting their applicability in unconstrained, real-world settings. Still, phase representations remain a valuable signal modality and, when combined with other learning-based methods, may help enhance robustness against photometric and structural distortions.
2.2.5 GAN-based
@@
A major limitation of conventional learning-based VFI approaches lies in their reliance on pixel-level loss functions such as $\ell_{1}$, $\ell_{2}$, or perceptual losses based on deep features (e.g., VGG [147]). While these objectives are effective for minimizing reconstruction errors, they often produce perceptually unsatisfying results, characterized by over-smoothed textures and diminished realism [148, 149]. To mitigate this gap, several methods adopt Generative Adversarial Networks (GANs) [150], which demonstrate remarkable performance in synthesizing visually plausible content [151, 152]. These GAN-based VFI methods [93, 101, 103, 74, 36, 94, 96, 95, 102, 97] employ a generator $G$ to synthesize the intermediate frame $\hat{I}_{t}$, and a discriminator $D$ to differentiate between the GT $I_{t}$ and $\hat{I}_{t}$. The generator is optimized using both a reconstruction loss and an adversarial loss, enabling it to produce frames that are structurally coherent with the inputs while exhibiting high perceptual fidelity. Such formulations are particularly effective in hallucinating plausible content in disoccluded regions [153] and enhancing visual details in blurry or textureless areas [74].
Despite their potential, GAN-based models introduce new challenges. They are notoriously difficult to train, often suffering from instability, mode collapse [154], and poor generalization when exposed to motion patterns or scene layouts not well represented in the training data. In such cases, the generator may fail to generalize, leading to artifacts or unrealistic interpolations. As a result, domain adaptation or fine-tuning is often required when applying GAN-based methods to novel environments [155], limiting their scalability in practical deployment.
2.2.6 Transformer-based
@@
Originally proposed for sequence modeling in natural language processing (NLP) [123], the Transformer architecture has been successfully adapted to VFI [50, 51, 52, 53, 54, 55, 56, 124, 113, 57, 117] owing to its strong capacity for capturing long-range dependencies through the attention mechanism [123, 156]. In the context of VFI, where motions often span large spatial and temporal regions with occlusions and deformations, this capability is particularly advantageous. The attention mechanism adaptively weighs features by their relevance to selectively attend to distant yet semantically relevant regions. This is an essential property for synthesizing temporally coherent intermediate frames. The core attention operation is defined as:
@@
where $Q$, $K$, and $V$ denote the query, key, and value matrices, respectively, and $d$ is the dimensionality of the feature space. This formulation enables the model to focus on spatial-temporal regions that are informative for interpolation, while effectively handling occlusions and appearance changes [37].
@@
Transformer-based VFI methods [50, 51, 52, 53, 54, 55, 56, 124, 113, 57, 117] primarily differ in how they structure attention and encode temporal dependencies. VFIFormer [52] introduces a cross-scale window-based attention (CSWA) mechanism to capture multi-scale dependencies without relying on flow-based motion estimation. Queries are computed from features at the target time, while keys and values are derived from neighboring input frames, enabling direct temporal associations. The multi-scale windowing expands the receptive field, enhancing robustness to complex motion. VFIT [51] employs a hierarchical Transformer operates on multi-resolution features and predicts spatially adaptive blending kernels for fine-grained synthesis. EMA-VFI [53] integrates attention modules with CNNs to reduce overhead, using inter-frame attention to jointly extract motion and appearance cues with improved efficiency.
@@
Despite their effectiveness, this approach suffers from high computational costs. The standard self-attention scales quadratically with the input resolution, posing a bottleneck for HR video inputs. To overcome this, efficient attention designs have been proposed. Swin Transformer [156] reduces complexity via windowed self-attention and shifted windows, while Restormer [157] introduces transposed attention to achieve linear complexity with respect to spatial dimensions. These developments point to a promising direction in which Transformer-based architectures may effectively balance global context modeling with computational efficiency, enabling real-time HR frame interpolation in practical applications.
2.2.7 Mamba-based
Structured State Space Models (SSMs) [98] offer a principled framework for sequence modeling through continuous-time dynamical systems. Among them, Mamba [99] introduces a selective state-space parameterization that combines the recurrent efficiency of recurrent neural networks (RNNs) [158] with the global context modeling capabilities of Transformers. By leveraging linear recurrence and input-dependent gating, Mamba enables long-range dependency modeling with linear complexity. The core of SSM-based modeling is the continuous-time linear time-invariant (LTI) system defined as:
$h^{\prime}(t)$ $=Ah(t)+Bx(t),$ (14)
$y(t)$ $=Ch(t)+Dx(t),$
where $h(t)\in\mathbb{R}^{N}$ is the latent state, $x(t)\in\mathbb{R}$ is the input, and $y(t)\in\mathbb{R}$ is the output. Here, the state size is $N$, with system parameters $A\in\mathbb{R}^{N\times N},\quad B\in\mathbb{R}^{N\times 1},\quad C\in\mathbb{R}^{1\times N},\quad D\in\mathbb{R}$. To incorporate this formulation into deep learning models, the system is typically discretized using the zero-order hold (ZOH) method with a step size $\Delta$. The resulting discrete parameters are computed as:
$\bar{A}$ $=\exp(\Delta A),$ (15)
$\bar{B}$ $=(\Delta A)^{-1}(\exp(\Delta A)-I)\Delta B,$
which yields the following discrete-time recurrence:
$h_{k}$ $=\bar{A}h_{k-1}+\bar{B}x_{k},$ (16)
$y_{k}$ $=Ch_{k}+Dx_{k}.$
which defines the fundamental update rule for state-space sequence modeling.
VFIMamba [100] is the first work to incorporate Mamba into VFI. It introduces a hierarchical architecture based on the S6-based Mixed-SSM Block (MSB) to model temporal dynamics across spatial resolutions. This design enables bidirectional propagation of motion features through structured recurrence, effectively capturing both short- and long-range dependencies. Compared to Transformer-based models [50, 51, 52, 53, 54, 55, 56, 124, 113, 57, 117], VFIMamba achieves lower memory consumption and faster inference while maintaining competitive accuracy, particularly in handling large displacements and preserving high-frequency texture details. MambaFlow [140] presents a Mamba-centric framework for end-to-end optical flow estimation. It extends the modeling capacity of Mamba through two core mechanisms. Self-Mamba captures long-range intra-frame dependencies by applying bidirectional state updates to enrich spatial features with global context. Cross-Mamba, inspired by cross-attention, models inter-frame interactions to improve motion correspondence. Together, these modules collectively improve robustness to occlusion, motion discontinuities, and ambiguous flow regions, making the architecture a promising backbone for VFI. Additionally, other SSM-based models such as MambaIR [159] and MambaIRv2 [160] demonstrate strong performance in image restoration tasks by capturing local detail and global structure with low complexity. The success of these models suggests that structured recurrence offers a compelling alternative to attention mechanisms for spatiotemporal modeling in VFI.
Overall, Mamba provides a promising design space for future VFI frameworks. Its low architectural complexity and reduced sensitivity to hyperparameter tuning offer practical advantages over attention-based designs. However, its ability to handle long-term motion dependencies, occlusion, and non-rigid deformation remains underexplored. Future directions may include exploring hybrid architectures that combine Mamba with local attention or design adaptive recurrence mechanisms conditioned on motion complexity and occlusion patterns.
2.3 Diffusion Model-based
@@
Figure 7: General structure of DM-based VFI framework. The framework receives input keyframes ($I_{0}$, $I_{1}$) and generates intermediate frames ($I_{t}$) through a denoising process. In addition to input keyframes, the model can accept various auxiliary conditioning signals such as images, text, audio, optical flow, or semantic maps via lightweight adapter modules or attention mechanisms.
Traditional deep learning-based VFI methods are predominantly deterministic, assuming a one-to-one mapping between the input frames and the interpolated output [113]. While such models demonstrate strong performance under moderate motion, they encounter fundamental limitations when faced with severe occlusions, significant displacements, or rapid appearance changes, where the underlying motion is ambiguous. In such cases, a deterministic framework cannot fully capture the range of plausible transitions between two frames, often leading to results that diverge from human perceptual expectations [161]. These limitations have prompted the exploration of a generative approach that embraces uncertainty and seeks to synthesize diverse yet semantically coherent interpolations.
DMs [162, 163, 164, 165, 166, 167] have emerged as a powerful generative model, achieving state-of-the-art (SOTA) performance across image [168, 164], video [165, 166], and multimodal generation [169, 170]. Unlike GANs [150] or VAEs [171], which suffer from adversarial instability and posterior collapse respectively, DMs offer stable training, high-fidelity samples, and strong temporal consistency. Inspired by their success in text-to-video (T2V)) [172, 173, 174] and image-to-video (I2V) [175, 169], researchers have recently adapted DMs for VFI [104, 105, 106, 107, 114, 116, 117], expanding the scope of VFI from deterministic interpolation to conditional generative modeling. This paradigm shift redefines VFI as a conditional denoising process, aligning it with the broader concept of Generative Inbetweening [112, 111], which focuses on synthesizing plausible and temporally coherent transitions between sparse keyframes under uncertainty. In this formulation, VFI is modeled as a denoising process conditioned on keyframes, typically denoted as $I_{0}$ and $I_{1}$. Starting from Gaussian noise, a denoising network gradually synthesizes intermediate frames via learned denoising steps. For example, Stable Video Diffusion (SVD) [167] adopts a latent diffusion framework. It first encodes video sequences into compact representations via an encoder $\mathcal{E}(\cdot)$, adds noise in the latent space, and then denoises them using a 3D U-Net [25]. The denoising objective, such as the $\mathbf{v}$-prediction loss [176], encourages accurate reconstruction from noisy inputs:
@@
where $\mathbf{v}=\alpha_{t}\epsilon-\sigma_{t}\mathbf{z}_{t}$, $\mathbf{z}_{t}$ is the noisy latent at timestep $t$, $\epsilon$ is the GT noise, and $\alpha_{t}$, $\sigma_{t}$ are variance schedule parameters that define the weighting between signal and noise. $\mathbf{c}_{\text{image}}$ denotes the conditioning keyframes.
Initial DM-based VFI models [104, 105] directly predict intermediate content without relying on explicit motion estimation. Building upon this foundation, more recent methods [114, 111] extend VFI into a semantically aware generation task, emphasizing coherent scene evolution over time. To enhance temporal alignment, TRF [112] and ViBiDSampler [115] propose bidirectional sampling and trajectory fusion, facilitating robust interpolation from both forward and backward perspectives without task-specific training. Beyond inference strategies, architectural innovations improve motion control and temporal consistency. EDEN [117] augments the denoising network with a spatio-temporal encoder for global consistency, while MoG [107] integrates motion priors using flow-guided warping in the latent space. As shown in Fig. 7, one of the key strengths of DM-based frameworks lies in their inherent flexibility to incorporate diverse conditioning modalities beyond input keyframes. DMs can seamlessly integrate auxiliary signals such as depth, semantic maps, audio, text, or motion priors via adapter-based [109, 108] or attention-based conditioning pathways. This allows for rich user guidance and semantic control, enabling use cases like such a story-driven animation [108], cross-modal interpolation, and interactive video generation [161]. Framer [161] injects spatial priors into the U-Net via attention mechanisms, while MoG [107] and FCVG [114] adopt ControlNet-like structures [177] to condition the generative process at multiple scales, improving alignment and spatial consistency.
Overall, DM-based VFI provides a new perspective on VFI by decoupling interpolation from deterministic regression and introducing generative modeling as a robust alternative. The ability of DMs to integrate diverse conditioning modalities and to model uncertainty enables flexible and perceptually plausible frame synthesis. However, their computational cost and sampling latency remain notable challenges. Future research directions may include hybrid frameworks that combine explicit motion estimation with generative denoising, as well as curriculum- or cascade-based denoising strategies tailored for HR inputs or temporally long-range interpolation tasks.
3 Learning Paradigm
3.1 Center-Time Frame Interpolation (CTFI)
@@
Figure 8: Comparison of CTFI and ATFI. (a) CTFI only generates a single center-frame at $t{=}0.5$ given two inputs. (b) ATFI can synthesize frames at arbitrary $t\in(0,1)$.

### v3

@@ II-B3 Kernel- and Flow-based Combined
Kernel- and flow-based approaches each offer distinct strengths in VFI. Flow-based methods explicitly model pixel-wise motion to enable temporally consistent frame alignment, but are sensitive to inaccuracies in optical flow estimation. In contrast, kernel-based methods directly synthesize pixels using learned, spatially adaptive convolutional kernels, where motion cues are encoded implicitly in the sampling support and kernel weights rather than in an explicit dense flow field, offering greater robustness in regions with complex motion. However, they are limited by their local receptive field and thus struggle with large displacements.
From a motion-representation standpoint, hybrid designs make explicit that kernel- and flow-based strategies are complementary rather than mutually exclusive. Flow fields provide an explicit description of pixel-wise displacement that is easy to visualize, debug, and reuse for arbitrary timestamps, whereas kernel-based sampling offers an implicit representation in which motion is captured by content-adaptive sampling locations and aggregation weights, often yielding more stable behavior in low-texture or motion-ambiguous regions. Hybrid architectures exploit this complementarity by using flow to provide globally coherent displacement guidance, while kernels refine local appearance and compensate for residual misalignment.
@@
Hybrid methods combine these advantages by leveraging optical flow to guide the placement and orientation of learned convolutional kernels, achieving global motion alignment while enabling local refinement. This combined approach [167, 6, 7, 262, 177, 124, 171, 46] typically begins with estimating optical flows using dedicated or pre-trained flow networks [108, 236, 56, 99, 186, 212, 96, 8, 218, 91, 250], which provide the sampling offsets for adaptive kernels. These kernels are then applied along flow-guided paths to aggregate motion-aware pixel neighborhoods. MEMC-Net [7], for example, combines PWC-Net [212] for flow estimation and deformable convolution [40] for localized refinement. In this setup, flow fields define the sampling offsets, while the kernel weights are learned to capture residual motion and restore high-frequency content.
@@ II-B4 Phase-based
An alternative direction in VFI exploits the phase information to capture motion cues. In the frequency domain, pixel-wise representations can be decomposed into amplitude and phase components, where temporal phase shifts across frames encode the apparent motion of underlying structures. To extract and manipulate phase information, most phase-based methods [159, 158] adopt multi-scale frequency representations such as complex steerable pyramids [205, 206, 185]. Motion is then modeled by interpolating both phase and amplitude at each pyramid level. Meyer et al[159] solves this optimization problem explicitly, while PhaseNet [158] adopts end-to-end learning strategies. These methods offer robustness to lighting changes and subpixel motion, without relying on explicit pixel correspondence.
@@
However, the underlying assumption that motion can be approximated as local phase shift, fails under large displacement. It leads to phase ambiguity and aliasing artifacts [227, 50]. As a result, phase-based methods often struggle to handle high-speed motion and often produce blurry results around sharp edges or occlusion boundaries. Nonetheless, phase representations remain a valuable signal modality and, when combined with other learning-based methods, may help enhance robustness against photometric and structural distortions.
II-B5 GAN-based
@@
Conventional learning-based VFI approaches predominantly rely on pixel-wise losses such as $\ell_{1}$, $\ell_{2}$, or deep feature-based perceptual losses (e.g., VGG [143]). Although these objectives effectively reduce reconstruction errors, they often lead to over-smoothed textures and lack of fine details, thereby compromising perceptual realism [156, 43]. To overcome this limitation, several methods adopt Generative Adversarial Networks (GANs) [65], which demonstrate remarkable performance in synthesizing visually plausible content [69, 11]. GAN-based VFI methods [224, 237, 262, 124, 46] employ a generator $G$ to synthesize the intermediate frame $\hat{I}_{t}$, while a discriminator $D$ distinguishes between the GT $I_{t}$ and $\hat{I}_{t}$. The generator is trained with both reconstruction and adversarial losses, enabling it to preserve structural consistency with the input frames while enhancing visual fidelity. Such formulations are particularly effective in hallucinating plausible textures in disoccluded or low-texture regions [123, 262].
Despite their potential, this approach introduces new challenges, including training instability, mode collapse [3], and limited generalization to unseen motion dynamics or scene layouts. These issues can lead to artifacts or unrealistic interpolations, especially when the training data lacks sufficient diversity. Consequently, domain adaptation or fine-tuning is often required when deploying these models in novel settings [25], raising concerns about scalability and robustness.
II-B6 Transformer-based
@@
Originally proposed for sequence modeling in natural language processing (NLP) [225], the Transformer architecture has been successfully adapted to VFI [36, 203, 150, 270, 133, 176, 266, 138, 152, 139, 279] owing to its strong capacity for capturing long-range dependencies through the attention mechanism [225, 147]. In the context of VFI, where motion often spans large spatial and temporal regions with occlusions and deformations, this ability is particularly advantageous. The attention mechanism adaptively weighs features by their relevance to selectively attend to distant yet semantically relevant regions. This is an essential property for synthesizing temporally coherent intermediate frames. The core attention operation is defined as:
@@
where $Q$, $K$, and $V$ denote the query, key, and value matrices respectively and $d$ is the dimension of the feature space. This formulation enables the model to focus on spatial-temporal regions that are informative for interpolation, while effectively handling occlusions and appearance changes [68].
@@
Transformer-based VFI methods primarily differ in how they structure attention and encode temporal dependencies. VFIFormer [150] introduces a cross-scale window-based attention (CSWA) mechanism to capture multi-scale dependencies without relying on flow-based motion estimation. Queries are computed from features at the target time, while keys and values are derived from neighboring input frames, enabling direct temporal associations. The multi-scale windowing expands the receptive field, enhancing robustness to complex motion. EMA-VFI [270] integrates attention modules with CNNs to reduce overhead, using inter-frame attention to jointly extract motion and appearance cues with improved efficiency.
@@
Despite their advantages, this approach faces high computational burden because standard self-attention scales quadratically with input size. To overcome this, efficient attention designs have been proposed. Swin Transformer [147] reduces complexity via windowed self-attention and shifted windows, while Restormer [265] introduces transposed attention to achieve linear complexity with respect to spatial dimensions. These developments point to a promising direction in which Transformer-based architectures may effectively balance global context modeling with computational efficiency, enabling real-time HR frame interpolation in practical applications.
II-B7 Mamba-based
Structured State Space Models (SSMs) [67] offer a principled approach to sequence modeling by formulating input–output dynamics through linear dynamical systems. Mamba [66] introduces selective state-space parameterization with input-dependent gating and linear recurrence, enabling efficient long-range dependency modeling with linear time complexity. This architectural simplicity and scalability make Mamba a compelling alternative to conventional attention-based models, particularly in scenarios where both temporal context length and computational budget are critical.
VFIMamba [267] is the first to adopt Mamba as a core temporal modeling backbone. Its hierarchical SSM-based design allows bidirectional recurrence across multiple spatial scales, facilitating robust motion feature propagation and long-range temporal alignment while maintaining low memory usage. In this sense, VFIMamba is representative of a new line of lightweight VFI architectures that replace global attention with structured recurrence, achieving competitive interpolation accuracy under substantially reduced computational cost. LC-Mamba [105] further refines this idea by incorporating shifted-window mechanisms and Hilbert-curve-based spatial scanning to preserve locality and continuity, which are crucial for high-resolution frame synthesis. These designs highlight the capacity of structured recurrence to capture both global motion trends and fine-grained dynamics necessary for high-fidelity interpolation. Beyond VFI, Mamba-based models such as MambaIR [71] and MambaIRv2 [70] have shown promising results in image restoration, suggesting that the core modeling principles behind Mamba generalize well across vision domains and tasks.
Taken together, these developments suggest that Mamba is emerging as a promising backbone for spatio-temporal modeling in VFI, particularly when lightweight deployment and long-range temporal context are required. At the same time, its behavior under challenging conditions such as severe occlusion or highly non-rigid motion remains relatively unexplored. Potential directions include localized or deformable recurrence, motion-aware conditioning, and hybrid designs that combine Mamba with complementary modules (e.g., occlusion-aware or non-linear-motion modeling components) to better handle complex video dynamics. We anticipate that future Mamba-based VFI architectures will increasingly exploit such combinations to improve expressiveness while preserving the favorable efficiency properties of structured state-space models.
II-C Diffusion Model-based
Diffusion Models (DMs) [82, 49, 190] have become a dominant generative framework in image [49, 190], video [83, 15], and multimodal synthesis [9, 274]. Compared to GANs [65] and VAEs [119], DMs offer more stable training, higher fidelity outputs, and better temporal coherence. Following their success in text-to-video (T2V) [242, 258, 263] and image-to-video generation (I2V) [189], recent research has extended DMs to VFI [231, 48, 268, 217, 226, 47, 102, 94, 287, 279, 228, 98, 255, 151, 27, 127, 74, 84].
@@ II-C Diffusion Model-based
Figure 7: General structure of DM-based VFI. The model receives input keyframes ($I_{0}$, $I_{1}$) and generates intermediate frames ($I_{t}$) through a denoising process. In addition to input keyframes, the model can accept various auxiliary conditioning signals such as images, text, audio, optical flow, or semantic maps via adapter modules or attention mechanisms.
Among diffusion backbones, Stable Video Diffusion (SVD) [14] first encodes video frames into a latent space via an encoder $\mathcal{E}(\cdot)$, adds Gaussian noise, and then denoises the representation using a 3D U-Net [39]. A typical loss formulation is the $\mathbf{v}$-prediction objective [192]:
@@
where $\mathbf{v}=\alpha_{t}\epsilon-\sigma_{t}\mathbf{z}_{t}$, $\mathbf{z}_{t}$ is the noisy latent at timestep $t$, $\epsilon$ is the GT noise, and $\alpha_{t}$, $\sigma_{t}$ are variance schedule parameters that define the weighting between signal and noise. $\mathbf{c}_{\text{image}}$ denotes input frames as condition. This reframes VFI as a conditional generation task in latent space.
Early DM-based VFI works such as MCVD [226] and LDMVFI [47] generate intermediate frames directly from noise conditioned on keyframes, without explicitly modeling motion. As shown in Fig. 7, a key strength of DMs is their ability to support flexible conditioning on auxiliary signals such as optical flow, semantic maps, audio, or text through adapters [272, 183, 87] or cross-attention mechanisms. For example, MoG [94] and FCVG [287] employ ControlNet [272] to inject motion priors into the denoising process, while Framer [231] integrates spatial priors via attention-based guidance.
Recently, VFI has been generalized into a broader task termed Generative Inbetweening [59, 232, 84]. Unlike conventional VFI, which assumes short temporal gaps between similar input frames, this formulation handles sparse and semantically distant input keyframes. However, this new concept introduces greater motion ambiguity, making temporal alignment more challenging. To address this, bidirectional strategies like TRF [59] and ViBiDSampler [254] fuse forward and backward sampling trajectories. On the architectural side, EDEN [279] employs a spatiotemporal encoder to enhance global consistency, while TLB-VFI [151] utilizes 3D-wavelet gating and temporal-aware autoencoding for motion fidelity. Another recent development is the emergence of Diffusion Transformers (DiTs) [181]-based VFI methods [84, 268]. ArbInterp [268] further emphasizes temporally flexible sampling, enabling generation at arbitrary timestamps with content-aware temporal control. Together with TRF [59] and ViBiDSampler [254], it exemplifies diffusion-guided temporal modeling in which the sampling trajectory is regularized by motion-aware priors rather than being purely stochastic, leading to improved global coherence over long sequences.
In parallel, cross-modal and multi-modal conditioning has been explored as a means to reduce motion ambiguity. MultiCOIN [217] performs controllable video inbetweening by injecting textual, visual, and structural controls (e.g., trajectories or layout cues) that jointly steer geometry and appearance, while BBF [48] leverages audio–visual semantic guidance to improve context-aware interpolation around boundary frames, indicating that audio cues can help disambiguate motion direction and event timing in complex scenes.
Despite their potential, DM-based approach faces several limitations: high computational cost, slow sampling, and limited scalability to long-range or HR settings. To mitigate these issues, lightweight or training-free designs [255, 228] have been explored, along with frameworks that decouple motion estimation from generative synthesis [74, 127], offering improved efficiency and controllability.
Overall, DM-based VFI has progressed from straightforward conditional denoising between nearby frames to generative inbetweening frameworks that incorporate diffusion-guided temporal modeling, multi-modal conditioning, and increasingly efficient architectures. We anticipate that DMs will play an important role in VFI, particularly in scenarios involving sparse keyframes, ambiguous motion, and high-level user controls, while further advances in efficiency and controllability will be crucial for practical deployment.
@@
Figure 8: Comparison of CTFI and ATFI. (a) CTFI only generates a single center-frame at t=0.5t{=}0.5 given two inputs. (b) ATFI can synthesize frames at arbitrary $t\in(0,1)$.
III Learning Paradigm
III-A Center-Time Frame Interpolation (CTFI)